### v1

# Conveyor-belt superconducting quantum computer

Abstract
The processing unit of a solid-state quantum computer consists in an array of coupled qubits, each locally driven with on-chip microwave lines that route carefully-engineered control signals to the qubits in order to perform logical operations. This approach to quantum computing comes with two major problems. On the one hand, it greatly hampers scalability towards fault-tolerant quantum computers, which are estimated to need a number of qubits – and, therefore driving lines – on the order of $10^{6}$. On the other hand, these lines are a source of electromagnetic noise, exacerbating frequency crowding and crosstalk, while also contributing to power dissipation inside the dilution fridge. We here tackle these two overwhelming challenges by presenting a novel quantum processing unit (QPU) for a universal quantum computer which is globally (rather than locally) driven. Our QPU relies on a string of superconducting qubits with always-on ZZ interactions, enclosed into a closed geometry, which we dub “conveyor belt”. Strikingly, this architecture requires only $\mathcal{O}(N)$ physical qubits to run a computation on $N$ computational qubits, in contrast to previous $\mathcal{O}(N^{2})$ proposals for global quantum computation. Additionally, universality is achieved via the implementation of single-qubit gates and a one-shot Toffoli gate. The ability to perform multi-qubit operations in a single step could vastly improve the fidelity and execution time of many algorithms.
I Introduction
Scalability in Quantum Computing (QC) refers to the ability to increase the number of qubits and other resources in a quantum system while maintaining performance, error rates, and coherence times Fowler2012 . Achieving scalability is essential for building quantum computers capable of solving complex problems beyond the reach of classical computers. For devices based on superconducting circuits Gambetta2017 ; Martinis2019quantum ; SQ2020 ; supremacy_PRL_2021 ; blais_review_RMP ; bravyi2022future ; ezratty2023perspective (and other solid-state platforms), scalability is fundamentally constrained, among others, by the need for localized drive of each physical qubit You2005supercond ; Girvin2008 ; devoret2013supercond ; wedin2017quantum in order to perform single-qubit rotations and entangling operations as required by the computation. This issue, often called the “wiring problem”, leads to a wiring overload Gambetta2017 ; tsai2020pseudo ; tsai2021gate : for example, if we had to scale the current microwave interface to control a million-qubit system, the electronics itself would take up to three football fields of space, while consuming about 40 MW of dc power Bardin2020 . A potential solution to this problem is the development of globally driven QC schemes Lloyd_1993 ; benjamin_2000 ; benjamin_2001 ; benjamin_2003 ; benjamin_2004 . These schemes involve designing and operating quantum computers in a manner that minimizes or eliminates the need for precise, localized manipulation of individual qubits. The main benefits include the simplification of designs due to the reduced complexity of control electronics and interconnections, and potentially enhanced error resilience with respect to local errors. However, globally driven QC approaches also face challenges, such as ensuring precision in global operations, creating and maintaining large, possibly highly entangled states for logical information encoding, and implementing fault-tolerant operations globally, which require sophisticated error-correcting protocols.
A recently proposed globally driven superconducting QC architecture menta2024globally addresses some of these challenges, specifically the precision in global operations and effectiveness of state preparation. Building upon a proposal based on Rydberg atoms cesa2023universal ; fromonteil2024hamilton , in Ref. menta2024globally we demonstrated that a globally driven universal quantum computer can be achieved in a superconducting platform by leveraging the longitudinal (always-on) ZZ interaction ZZ-suppression2020 ; Xu2024 ; ZZ-contrast2020 ; ZZ-eliminating2022 ; ZZ_2020 ; ZZ_2021 ; xu2021z ; Fors2024 between neighbouring superconducting qubits organized in a two-dimensional (2D) ladder. In this work we present a novel architecture that significantly improves the design proposed in Ref. menta2024globally by drastically reducing the resources needed to implement it.
Figure 1: Schematic description of the proposed architecture. Two types, $A$ and $B$, of superconducting qubits (red and blue squares, respectively) are separately driven by two classical sources $V_{A,B}(t)$ (red and blue continuous lines). They are coupled via a longitudinal ZZ coupling (black and grey springs). Black triangles inside regular and crossed qubits denote local corrections in the resonance frequency of the qubits (see main text). The $A$-type crossed qubit (red square inside the loop) enables one-shot Toffoli gate (three-qubit gate) – the corresponding interactions are depicted in gray. The $B$-type crossed qubit performs single-qubit gates. The elements highlighted in yellow indicate the information carrying sites $Q_{1}$, $Q_{2}$, $\cdots$, $Q_{N}$, separated by three-qubits sectors $S_{1}$, $S_{2}$, $\cdots$, $S_{N}$. The $Q_{j}$’s host the computational qubits through well-formed configurations (2) where the $S_{j}$’s are initialized in alternating sequences of paramagnetic ($geg$) and ferromagnetic ($ggg$) phases. Since ZZ interactions critically depend on the relative distance between two qubits, we have used the honeycomb structure to make sure that the couplings among adjacent qubits are (nominally) all identical. The figure pertains to a $N=8$ qubit quantum computer.
At variance with previous schemes menta2024globally ; cesa2023universal , which require to operate globally on a 2D array of ${\cal O}(N^{2})$ physical qubits to run universal operations on $N$ computational qubits, the setup discussed here only needs ${\cal O}(N)$ elements – specifically $14N+$ 1 – organized in a closed loop, as shown in Fig. 1. For clarity, we stress that the computational qubits are the ones in which we encode the logical information. On the other hand, the set of physical qubits encompasses all two-level systems present in the architecture. Sequences of control pulses that act collectively on half of the elements of the device allow us to induce special permutations on the states of the sites. Combining these transformations with the possibility to implement both single-qubit gates and a one-shot Toffoli gate yields universal quantum computation. It is important to mention that the associated dynamical processes are based on a blockade regime, analogous to the one in Rydberg atoms Rydberg_blockade , which is here achieved thanks to a strong ZZ coupling between nearest-neighbor qubits. The main features of this effect for the case of superconducting qubits are described in Ref. menta2024globally . We also mention that, in QC, the ability to perform multi-qubit operations, such as the Toffoli gate, in a single step could vastly improve the fidelity and execution time of many algorithms Zahedinejad2015 ; baker2022 . To the best of our knowledge, our proposal introduces a novel superconducting platform that, for the first time, incorporates a one-shot Toffoli gate Nielsen2010 , which operates based on the ZZ-induced blockade regime.
II Results
II.1 Model
The system shown in Fig. 1 consists of a closed loop formed by $4N$ superconducting qubits coupled via a spatially-uniform ZZ interaction of coupling strength $\hbar\zeta$ (black springs in the figure). The loop contains $N$ sites, identified by the symbols $Q_{1}$, $Q_{2}$, $\cdots$, $Q_{N}$, which play a special role in the model and will be referred to as Information Carrying (IC) sites. For all $j\in\{1,\cdots,N\}$ the IC site $Q_{j}$ is separated from its adjacent counterpart $Q_{j+1}$ by a sector $S_{j}$ which contains three ”non-active” sites. In the figure, there is an extra superconducting qubit placed inside the loop, which is connected to the first three IC sites $Q_{1}$, $Q_{2}$, and $Q_{3}$ via the same ZZ coupling term that connects every other element (these three couplings are highlighted in a light gray color to emphasize the fact that this additional qubit is placed inside the loop and not along it). The superconducting qubits in the model belong to two distinct families: $A$-type qubits, represented as red squares, and $B$-type qubits, represented as blue squares, which alternate in the loop in an $ABA$ pattern. As shown in the figure, all the IC sites host a $B$-type qubit, while the intermediate sectors $S_{j}$ are formed by one $B$-type and two $A$-type qubits. All qubits of the same type share the same level spacing $\hbar\omega_{A/B}$, except for the elements marked with triangles, which, for reasons that will be clear in the following, require $\zeta$-dependent local corrections to their level spacings i.e. $\hbar\omega_{A/B}\to\hbar(\omega_{A/B}+\zeta)$, due to the higher number of nearest neighbors they have (three) compared to the other qubits in the setup. To eliminate the possibility of a residual “swap” term between nearest-neighbor qubits, it may be necessary to have $\omega_{A}\neq\omega_{B}$. Each family of qubits is collectively driven by the same time-dependent external source, termed $V_{A/B}(t)$. Thus, the control is global. The device also includes two inhomogeneities: the additional $A$-type element previously mentioned and one of the three $B$-type elements connected to it (specifically, the one located at the IC site $Q_{2}$). These special qubits, referred to as crossed qubits, are used to perform multi- and single-qubit operations. They maintain the nominal level spacing of their respective families but have an augmented coupling with the external source, i.e. the Rabi frequency. As shown in Ref. menta2024globally , this difference allows independent control of the normal qubits and crossed qubits of a given family, despite the global control pulse acting simultaneously on them. Both the $A$- and $B$-type crossed qubits are marked with a small black triangle. Finally, a third control line, $V_{\mathrm{init}}(t)$, acting on a subset of the $B$-type qubits, is necessary for the initialization of the machine. This control line is active only at the initial stage of the computation. Regarding the readout procedure, a possible implementation is discussed in the supplemental material supplemental .
Adopting the same convention of Ref. menta2024globally we write the Hamiltonian of the setup as $\hat{H}(t):=\hat{H}_{0}+\hat{H}_{\rm drive}(t)$, where $\hat{H}_{0}:=\sum_{\chi\in\{A,B\}}\sum_{i\in\chi}\frac{\hbar\omega_{i}}{2}\hat{\sigma}_{i}^{(z)}+\sum_{\langle i,j\rangle}\frac{\hbar\zeta}{2}\hat{\sigma}_{i}^{(z)}\otimes\hat{\sigma}_{j}^{(z)}$, describes the local energy contribution of the superconducting qubits and their ZZ interactions which are fixed by the geometry of the model, while $\hat{H}_{\rm drive}(t):=\sum_{\chi\in\{A,B\}}\sum_{i\in\chi}\hbar\Omega_{\chi}(t)\sin(\omega_{\mathrm{d},\chi}t+\phi_{\chi}(t))\hat{\sigma}_{i}^{(y)}$, is the time-dependent driving contribution induced by the classical control lines. In these equations, $\hat{\sigma}_{i}^{(x,y,z)}$ represent ordinary Pauli matrices acting on the Hilbert space of the $i$-th qubit, expressed in the local energy basis $\{|g_{i}\rangle,|e_{i}\rangle\}$. The summation in the interacting part of $\hat{H}_{0}$ encompasses all nearest-neighbor interactions, mathematically representing all the black springs in Fig. 1. The parameter $\omega_{\mathrm{d},\chi}$ denotes the oscillation frequency of the driving pulse $V_{\chi}(t)$, while $\Omega_{\chi}(t)$ and $\phi_{\chi}(t)$ define the time-dependent Rabi frequency and phase of such control. In the following, we will assume that these quantities assume constant values on disjoint time windows, in such a way that, at each time, only one species (either $A$ or $B$) is driven. For simplicity, in writing $\hat{H}_{\rm drive}(t)$, we have omitted the driving term associated with the control line $V_{\rm init}(t)$, which operates solely at the very beginning of the computational process on a specific subset of the physical qubits. Additionally, we did not explicitly highlight that when the site index $i$ in the expression for $\hat{H}_{0}$ identifies a $B$-type qubit marked with a black triangle, the corresponding qubit’s level spacing becomes $\hbar(\omega_{B}+\zeta)$ instead of $\hbar\omega_{B}$. The same substitution must be performed when the site index $i$ identifies an $A$-type qubit marked with a black triangle, i.e. an $A$-type crossed qubit. Similarly, whenever in the expression of $\hat{H}_{\rm drive}(t)$ the index $i$ identifies a crossed qubit, $\Omega_{\chi}(t)$ needs to be replaced by $2\Omega_{\chi}(t)$. Apart from these local adjustments (which need to be engineered once and for all at fabrication level), it is important to note that $\Omega_{\chi}(t)$, $\phi_{\chi}(t)$, and $\omega_{\mathrm{d},\chi}$ are independent of the site index $i$, indicating that they are associated with a control pulse acting globally on all qubits of $\chi$-type in the model. We conclude this Section by stressing that the above Hamiltonian is exactly the same as the one studied in Ref. menta2024globally . However, due to the more exotic geometry, the dynamical features of the system are more complex than in the case of the 2D ladder geometry used in Ref. menta2024globally . We take advantage of this complexity to achieve the $\mathcal{O}(N)$ scaling, as opposed to the $\mathcal{O}(N^{2})$ scaling of Refs. menta2024globally ; cesa2023universal . In the next paragraph we describe such features. Further details can be found in the supplemental material supplemental as well as in Ref. menta2024globally .
II.2 Dynamical features
The first ingredient of the present proposal is to emulate the Rydberg blockade effect Rydberg_blockade using the ZZ interactions of the model. Specifically, one can show that under the strong-coupling condition $\eta_{\rm BR}:=|\zeta/\Omega_{\chi}|\gg 1$, by properly detuning the driving frequencies $\omega_{\mathrm{d},\chi}$ from the nominal level spacing of the qubits, one is able to selectively induce transitions among states of the system only when two nearest-neighbour sites do not simultaneously occupy their excited levels. The second fundamental ingredient is the presence of two different values of the Rabi frequency for qubits of the same family, i.e. the presence of regular and crossed qubits. Crucially, one can prove that this difference is sufficient to drive independently these two types of qubits even though their control is global menta2024globally . Specifically, one is able to perform generic unitaries defined as
$\hat{W}_{\chi}(\theta^{\prime},\bm{n}^{\prime};\theta^{\prime\prime},\bm{n}^{\prime\prime}):=\hat{W}_{\chi^{\rm r}}(\theta^{\prime},\bm{n}^{\prime})\hat{W}_{\chi^{\times}}(\theta^{\prime\prime},\bm{n}^{\prime\prime})~{},$ (1)
[…]
II.3 Information encoding and exchange operations
In the setup of Fig. 1 the logical information is encoded in the IC sites $Q_{1}$, $Q_{2}$, $\cdots$, $Q_{N}$ with the intermediate sectors $S_{1}$, $S_{2}$, $\cdots$, $S_{N}$ acting as separators. This is a peculiarity of the present quantum computing processor, which has no analogue in the proposals of Refs. menta2024globally ; cesa2023universal , where the string of the $N$ computational qubits can instead rigidly drift along the entire 2D array of physical qubits of the device. Specifically, a generic $N$-qubit logical state $|\Psi\rangle=\sum_{\vec{k}\in\{e,g\}^{N}}\Psi_{\vec{k}}|k_{1},k_{2},\cdots,k_{N}\rangle$ is expressed in one of the two possible well-formed configurations, $|\Psi;{\rm FP}\rangle$ or $|\Psi;{\rm PF}\rangle$. Both these vectors have the central crossed $A$-type qubit in the ground state, while the intermediate sectors $S_{j}$ are in an alternating sequence of “ferromagnetic” ($|{\rm F}\rangle:=|ggg\rangle$) or “paramagnetic” ($|{\rm P}\rangle:=|geg\rangle$) phases. In particular, $|\Psi;{\rm FP}\rangle$ (resp. $|\Psi;{\rm PF}\rangle$) initializes the sector $S_{j}$ in the state $|{\rm F}\rangle_{S_{j}}$ if $j$ is odd (even), and in $|{\rm P}\rangle_{S_{j}}$ if $j$ is even (odd), so that
$|\Psi;{\rm FP}\rangle$ $:=\!\!\!\!\!\!\sum_{\vec{k}\in\{g,e\}^{N}}\Psi_{\vec{k}}|k_{1}\rangle_{Q_{1}}|{\rm F}\rangle_{S_{1}}|k_{2}\rangle_{Q_{2}}|{\rm P}\rangle_{S_{2}}|k_{3}\rangle_{Q_{3}}|{\rm F}\rangle_{S_{3}}$ (2)
$\cdots$ $|k_{N-1}\rangle_{Q_{N-1}}|{\rm F}\rangle_{S_{N-1}}|k_{N}\rangle_{Q_{N}}|{\rm P}\rangle_{S_{N}}\otimes|g\rangle_{A^{\times}},$
and $|\Psi;{\rm PF}\rangle=|\Psi;{\rm FP}\rangle|_{|{\rm P}\rangle\leftrightarrow|{\rm F}\rangle}$. We notice that this requires $N$ to be an even number. When all qubits are initially in the ground state, the control $V_{\rm init}(t)$ can be used to initialize the system to a vector of the form $|\Psi;{\rm FP}\rangle$, see Fig. 1. This process brings all the associated qubits into the $|e\rangle$ state. As mentioned before, in our architecture, single and multi-qubit gates can then be performed on specific sites which host, or which are directly coupled to, crossed qubits elements (i.e. $Q_{1}$, $Q_{2}$ and $Q_{3}$). This implies that a fundamental prerequisite to perform quantum computation in our conveyor-belt architecture is the ability to coherently exchange the positions of the computational qubits, i.e. change the positions of the qubit states without modifying them. This is achieved trough a sequence of eight alternating global pulses, i.e. $\hat{\Pi}_{\rm exc}:=\hat{\Pi}_{B}\hat{\Pi}_{A^{\rm r}}\hat{\Pi}_{B}\hat{\Pi}_{A^{\rm r}}\hat{\Pi}_{B}\hat{\Pi}_{A^{\rm r}}\hat{\Pi}_{B}\hat{\Pi}_{A^{\rm r}}$ with $\hat{\Pi}_{A^{\rm r}}:=\hat{W}_{A^{\rm r}}(\pi,\bm{x})$ acting as a conditional-bit-flip only on the regular $A$-type qubits, and $\hat{\Pi}_{B}:=\hat{W}_{B}(\pi,\bm{x};\pi,\bm{x})$ acting as a conditional bit-flip on all $B$-type qubits, including the crossed one. As shown in the supplemental material supplemental , when acting on a ${\rm FP}$ (resp. ${\rm PF}$) well-formed state $|\Psi;{\rm FP}\rangle$ ($|\Psi;{\rm PF}\rangle$), the transformation $\hat{\Pi}_{\rm exc}$ runs in parallel $N$ two-qubit swap gates $\hat{U}^{{\rm swap}}_{Q_{j}Q_{j+1}}$ on the pairs $\{Q_{1},Q_{2}\}$, $\{Q_{3},Q_{4}\}$, $\cdots$, $\{Q_{N-1},Q_{N}\}$, (resp. $\{Q_{2},Q_{3}\}$, $\{Q_{4},Q_{5}\}$, $\cdots$, $\{Q_{N},Q_{1}\}$) while exchanging the ferromagnetic and paramagnetic phases of the intermediate sectors producing a ${\rm PF}$ (resp. ${\rm FP}$) output configuration—see Fig. 2. The resulting motion of the computational qubits acquires an interesting feature. Indeed one has that $\hat{\Pi}_{\rm exc}|\Psi;{\rm FP}\rangle=|\Psi^{(1)}_{{\circlearrowright}_{o},{\circlearrowleft}_{e}};{\rm PF}\rangle$, and $\hat{\Pi}_{\rm exc}|\Psi;{\rm PF}\rangle=|\Psi^{(1)}_{{\circlearrowleft}_{o},{\circlearrowright}_{e}};{\rm FP}\rangle$, with the vector $|\Psi^{(\ell)}_{{\circlearrowright}_{o},{\circlearrowleft}_{e}}\rangle$ (resp. $|\Psi^{(\ell)}_{{\circlearrowleft}_{o},{\circlearrowright}_{e}}\rangle$) obtained by applying to $|\Psi\rangle$ an $\ell$-step clock-wise rotation of the internal states of the sites $Q_{j}$ with odd (even) index $j$ and, at the same time, an $\ell$-step anti-clock-wise rotation of the internal states of the sites $Q_{j}$ with even (odd) index $j$. Notice that since the direction of the rotations depends on the initial location of the ferromagnetic and paramagnetic phases supplemental , multiple applications of $\hat{\Pi}_{\rm exc}$ pulses do not cancel out. For instance, using $\ell$ times the transformation on the input state $|\Psi;{\rm FP}\rangle$ leads to $\hat{\Pi}^{\ell}_{\rm exc}|\Psi;{\rm FP}\rangle=|\Psi^{(\ell)}_{{\circlearrowright}_{o},{\circlearrowleft}_{e}};{\rm FP}\rangle$ for $\ell$ even, and $\hat{\Pi}^{\ell}_{\rm exc}|\Psi;{\rm FP}\rangle=|\Psi^{(\ell)}_{{\circlearrowright}_{o},{\circlearrowleft}_{e}};{\rm PF}\rangle$ for $\ell$ odd (of course for $\ell=N$ the system goes back to the initial configuration).
Figure 2: (a) Schematic description of the unitary evolution induced by $\hat{\Pi}_{\rm exc}$ on a ${\rm FP}$ well-formed state where the $Q_{7}$ and $Q_{8}$ sites are in the input states $|\psi_{7,8}\rangle$. This unitary operation acts as a swap between two qubits, separated by a ferromagnetic region. Moreover, after the unitary, paramagnetic and ferromagnetic phases are exchanged. The reverse unitary transformation is $\hat{\Pi}^{-1}_{\rm exc}:=\hat{\Pi}_{B}\hat{\Pi}_{\rm exc}\hat{\Pi}_{B}$. (b) Schematic representation of the universality proof.
Exploiting this feature, we can move any IC qubit in any other IC qubit location. Suppose, now, that the aim is to apply the single qubit transformation ^$\hat{\mathbb{R}}(\theta,\bm{n})$ to the computational qubit located (say) in the $Q_{j}$ site of the loop. Given our previous explanations, this task is straightforward. We simply need to use the transformation $\hat{\Pi}^{\ell}_{\rm exc}$ with $\ell$ such that the state of $Q_{j}$ rotates into the position $Q_{2}$ which is hosting the $B$-type crossed qubit. Once there, we apply the pulse $\hat{U}_{B^{\times}}:=\hat{W}_{B^{\times}}(\theta,\bm{n})$ and (in case we need it), reverse the exchange operation via the sequence $\hat{\Pi}^{N-\ell}_{\rm exc}$ (or using $\ell$-times the inverse of $\hat{\Pi}^{\ell}_{\rm exc}$ defined in the caption of Fig. 2). This demonstrates that we are capable of performing any single-qubit unitary operation on any one of the computational qubits. In the next Section, we describe how a three-qubit Toffoli gate can be implemented, thereby achieving a universal gate set.
II.4 One-shot Toffoli gate
In common QC architectures, the Toffoli gate is performed using two-qubit gates. Here, we implement a three-qubit one-shot Toffoli gate by using a total number of four qubits: one $A$-type crossed qubit located inside the loop, which acts as a mediator, and the three $B$-type computational qubits located in the IC sites $Q_{1}$, $Q_{2}$ and $Q_{3}$. A similar setup has already been proposed by Rasmussen et alRasmussen202_toffoli as an isolated multi-qubit gate. The most notable feature of our Toffoli gate is that it is naturally embedded into our conveyor-belt QC architecture. Here, the natural way to implement a Toffoli gate is to decompose it into a controlled-controlled-Z (CCZ) gate combined with two single-qubit Hadamard gates. Since we have already discussed how to perform single-qubit gates, we simply need to explain now how to perform a CCZ gate. The fundamental observation here is that the unitary $\hat{\mathcal{Z}}_{A^{\times}}:=\hat{W}_{A^{\times}}(2\pi,\bm{n})$ induces a $(-1)$ phase factor on the state of the system if and only if the three $B$-type qubits connected to it are all in the ground state $|g\rangle$ which, apart from a global NOT, is exactly a CCZ gate applied on the three $B$-type qubits. Specifically, the transformation $\hat{\mathcal{Z}}_{A^{\times}}$ is obtained from (1) for $\chi=A$ and by setting $\theta^{\prime}=0$ and $\theta^{\prime\prime}=2\pi$, independently from the choices of $\bm{n}^{\prime}$ and $\bm{n}^{\prime\prime}$. It corresponds to $\hat{\mathcal{Z}}_{A^{\times}}=\hat{Q}_{\langle A^{\times}\rangle}-\hat{P}_{\langle A^{\times}\rangle}$, where, since the $A$-type crossed qubit has three connections, the projector $\hat{P}_{\langle A^{\times}\rangle}$ can be written as $\hat{P}_{\langle A^{\times}\rangle}:=|ggg\rangle\langle ggg|$, while the complementary projector is $\hat{Q}_{\langle A^{\times}\rangle}:=|eee\rangle\langle eee|+|eeg\rangle\langle eeg|+\cdots+|gge\rangle\langle gge|$. Accordingly, setting $\bm{n}=(1,0,1)/\sqrt{2}$, the transformation
$\hat{T}_{13\rightarrow 2}:=\hat{W}_{B^{\times}}(\pi,\bm{n})\hat{W}_{B}(\pi,\bm{x})\hat{\mathcal{Z}}_{A^{\times}}\hat{W}_{B}(\pi,\bm{x})\hat{W}_{B^{\times}}(\pi,\bm{n})\ ,$ (3)
corresponds to a Toffoli gate where $Q_{1}$ and $Q_{3}$ act as controllers and $Q_{2}$ as the controlled qubit. In Eq. (3), $\hat{W}_{B}(\pi,\bm{x})$ stands for $\hat{W}_{B}(\pi,\bm{x};\pi,\bm{x})$.
We conclude this Section by remarking that this whole discussion can be generalized to an arbitrary number of qubits connected to the central $A$-type crossed qubit. It is therefore possible, at least in principle, to realize $N$-qubit Toffoli gates in our globally-controlled conveyor-belt architectures.
II.5 Universality
In the previous Sections, we have demonstrated how, by using the global controls $V_{A}(t)$ and $V_{B}(t)$, the architecture in Fig. 1 allows us to implement any single-qubit operation on each site that encodes the logical information, as well as to perform the exchange gate $\hat{\Pi}_{\rm exc}$ and the Toffoli gate $\hat{T}_{13\rightarrow 2}$. To prove that this is sufficient to guarantee universal QC in the model, we now show that, using these gates, it is possible to induce arbitrary swap gates $\hat{U}^{{\rm swap}}_{Q_{j}Q_{j^{\prime}}}$ between any two IC sites. These will then be used to convert $\hat{T}_{13\rightarrow 2}$ into Toffoli gates that couple all possible triples formed by computational qubits. It should first be noted that the Toffoli gate, when combined with single-qubit operations, can be employed to implement local swaps between any two qubits that are simultaneously connected to the central $A$-type crossed qubit Nielsen2010 . Starting from $\hat{T}_{13\rightarrow 2}$, we can generate swap gates between each of the pairs $\{Q_{1},Q_{2}\}$, $\{Q_{2},Q_{3}\}$, and $\{Q_{3},Q_{1}\}$ that we represent as edges of a $N$ vertex graph formed by the computational qubits of the system, see Fig. 2(b). Next, we apply $\hat{\Pi}_{\rm exc}$ to rotate the computational qubits, e.g. inducing the mapping $(Q_{1},Q_{2},Q_{3})\leftrightarrow(Q_{2},Q_{1},Q_{4})$, and repeat the whole procedure obtaining two new swaps among the couple $\{Q_{1},Q_{4}\}$ and $\{Q_{2},Q_{4}\}$ which allows us to draw two new edges in the graph. Note also that, by combining these swaps, we can also induce an extra new swap between $\{Q_{3},Q_{4}\}$ (e.g. $\hat{U}^{{\rm swap}}_{Q_{3}Q_{4}}=\hat{U}^{{\rm swap}}_{Q_{1}Q_{3}}\hat{U}^{{\rm swap}}_{Q_{1}Q_{4}}\hat{U}^{{\rm swap}}_{Q_{1}Q_{3}}$), making the subgraph formed by the sites $Q_{1}$, $Q_{2}$, $Q_{3}$ and $Q_{4}$ fully connected. Proceeding along this way, one can show by recursion that all the other vertices can also be included in a fully connected graph, meaning that we can generate all possible swaps $\hat{U}^{{\rm swap}}_{Q_{j}Q_{j^{\prime}}}$, hence proving the thesis supplemental .
III Discussion
The model we proposed represents a tremendous improvement in terms of scalability compared to those discussed in Refs. menta2024globally ; cesa2023universal where one needs $\mathcal{O}(N^{2})$ physical qubits in order to have a universal quantum computer with $N$ computational qubits. Indeed, in the present work, we have presented a quantum processing unit where the scaling is $\mathcal{O}(N)$. Globally-driven schemes with linear scaling have already been proposed  Lloyd_1993 ; benjamin_2000 ; benjamin_2001 ; benjamin_2003 ; benjamin_2004 . In contrast to our proposal, however, these ones remain rather abstract, since they do not provide precise recipes on how to actually implement the proposed global schemes. Additionally, there are many elements of difference in how the computation is performed. Among these, the processor described in the present work does not rely on the dynamical tuning of the system’s parameters, such as level spacings and couplings. Also, we do not use a redundant encoding of the computational qubits. A single, two-level system contains the information of a single computational qubit. This greatly simplifies in our humble opinion the extension of our conveyor-belt architecture to the level that is needed for fault-tolerant QC.
All considerations regarding the physical requirements (e.g. values of the Rabi frequencies, level spacings, ZZ coupling strengths, etc.) for the implementation of our setup remain consistent with those presented in Ref. menta2024globally , since the physical building blocks involved are identical. Another novelty introduced by the setup in Fig. 1, as compared to the one proposed in Ref. menta2024globally , is the ability of the former to perform one-shot Toffoli gates Zahedinejad2015 ; Rasmussen202_toffoli ; baker2022 . This could in principle improve the execution time as well as the fidelity of many quantum algorithms. From a practical standpoint, the fabrication of a single large loop, containing many qubits, and the design of global control lines that simultaneously connect all the qubits may prove to be an inefficient and error-prone process. Integrating a hybrid control with different global sources for different subsets of qubits, fabricated on different chiplets, could prove advantageous.
A major challenge of our architecture, from the experimental point of view, arises from the fact that spatial inhomogeneities in the fabrication of superconducting qubits and in the always-on ZZ couplings throughout the conveyor belt can influence the fidelity of the necessary logical operations. While a thorough disorder and noise analysis is well beyond the scope of this work, it is important to note that systematic fabrication errors can be managed as conventional dynamical errors within error correction frameworks.
To achieve fault-tolerant quantum computation, error-correction (EC) protocols are essential. As shown in Fig. 2(b), the number of computational qubits directly coupled to a given one is the same as in conventional 2D arrays, i.e. four connections. Therefore, in contrast with the proposal of Ref. menta2024globally , our system is equivalent to a 2D processor. This suggests that any standard fault-tolerant procedure can be implemented, e.g. surface codes Fowler2012 or qLDPC codes LDPC1 ; LDPC2 . Finally, although EC schemes for globally-controlled quantum systems have been explored previously Fitzsimons2009 , the development of globally-driven EC codes specific to our machine is a topic of great interest for future investigation.
We note in passing that, from an academic perspective, the setup depicted in Fig. 1 is, at least conceptually, similar to a “finite” quantum Turing machine Molina2019 ; QTM2020 . In this model, the $B$-type crossed qubit acts as a pointer affecting the quantum logical states of the machine, which we are able to move along the finite loop.
In conclusion, a similar quantum computing setup could be realized by substituting the Toffoli gate with two entangling gates. One of these should connect two even-indexed (or odd-indexed) computational qubits, while the other should connect two computational qubits of different parity. In this case, the universality of the computer would also be ensured. However, for it to be useful, one should be able to “activate” or “deactivate” the two $A$-type crossed qubits independently. Variants to the proposed setup that take into account this necessity are presented in Ref. supplemental . Finally, it should be noted that our setup can be implemented on different physical platforms, including Rydberg atoms Saffman_RMP_2010 and semiconductor spin qubits Review-Spin-qubits .
Acknowledgements.
It is a pleasure to thank M. Riccardi for useful comments and discussions.
Author Contributions
All authors contributed to developing the core idea of the paper and were equally involved in writing the manuscript.
Competing interests
M. P. and V. G. are co-founders of Planckian. R. A. is a scientist at Planckian and R. M. is a PhD scholar at Planckian. All other authors declare no competing interests.